OpenStudy (anonymous):

True or false. Teaching self-care skills is part of meeting the physical needs of toddler? PLEASE HELP

OpenStudy (anonymous):

https://www.youtube.com/watch?v=Fam3U_FGhR8

OpenStudy (anonymous):

I don't have any idea..

OpenStudy (anonymous):

\(KE = \dfrac{1}{2}mv^2\) Hitting the ground gives us KE = PE \(PE = mgh\) So we write: \(\dfrac{1}{2}mv^2 = mgh\) Multiply 2 to both sides: \(mv^2 = 2mgh\) Divide 'm' to both sides: \(v^2 = 2gh\) Now plug in 2100 and 4.50. \(v^2 = 2(2100)(4.50)\) What's 2 * 2100 * 4.50?

OpenStudy (anonymous):

Oh, wait actually we get: \(v = \sqrt{2(4.5)(9.8)}\) So what's 2 * 4.5 * 9.8? Can you mutliply those? @Moncee1995

OpenStudy (anonymous):

😭 I just want the answer! I want too finish my physics course ASAP

OpenStudy (anonymous):

It's easy, just multiply 2 * 4.5 * 9.8

OpenStudy (anonymous):

Just put it in your calculator

OpenStudy (anonymous):

On Open Study we have to guide askers to the answer, and we need their participation. So if you want the answer, you're gonna have to work for it.

OpenStudy (anonymous):

88.2?

OpenStudy (anonymous):

Yes, now we find the square root of 88.2 and we have our answer.